PROGRESS IN OPTICALLY-DETECTED SPIN-RESONANCE ON HELIUM DROPLETS

\underline{MARKUS KOCH}, GERALD AUBÖCK, JOHANN NAGL, CARLO CALLEGARI, and WOLF-GANG E. ERNST, Institute of Experimental Physics, TU Graz, Petersgasse 16, 8010 Graz, Austria/EU.

We have recently demonstrated the possibility of optically detecting the spin state of alkali metal atoms and molecules formed on helium nanodroplets and immersed in a magnetic field.^{a}^{b} This has already allowed us to show that the electronic spins of atoms do not relax within the timescale of the experiment (∼ 10^{−3} s) and those of molecules do. With this prerequisites fulfilled, we now set to demonstrate that electron spins on a He droplet can be manipulated. We have already achieved optical pumping, which we intend to present at the meeting, and we strive to soon reach optically-detected electron-spin resonance.

aJ. Nagl, G. Auböck, C. Callegari and W. E. Ernst, Phys. Rev. Lett. 98, 075301 (2007).

bG. Auböck, J. Nagl, C. Callegari and W. E. Ernst, J. Phys. Chem. A 111, 7404 (2007).